# PsychEval: A Multi-Session and Multi-Therapy Benchmark for High-Realism and Comprehensive AI Psychological Counselor

### v1

@@ Abstract
To develop a reliable AI for psychological assessment, we introduce PsychEval, a multi-session, multi-therapy, and highly realistic benchmark designed to address three key challenges: 1) Can we train a highly realistic AI counselor? Realistic counseling is a longitudinal task requiring sustained memory and dynamic goal tracking. We propose a multi-session benchmark (spanning 6-10 sessions across three distinct stages) that demands critical capabilities such as memory continuity, adaptive reasoning, and longitudinal planning. The dataset is annotated with extensive professional skills, comprising over 677 meta-skills and 4577 atomic skills. 2) How to train a multi-therapy AI counselor? While existing models often focus on a single therapy, complex cases frequently require flexible strategies among various therapies. We construct a diverse dataset covering five therapeutic modalities (Psychodynamic, Behaviorism, CBT, Humanistic Existentialist, and Postmodernist) alongside an integrative therapy with a unified three-stage clinical framework across six core psychological topics. 3) How to systematically evaluate an AI counselor? We establish a holistic evaluation framework with 18 therapy-specific and therapy-shared metrics across Client-Level and Counselor-Level dimensions. To support this, we also construct over 2,000 diverse client profiles. Extensive experimental analysis fully validates the superior quality and clinical fidelity of our dataset. Crucially, PsychEval transcends static benchmarking to serve as a high-fidelity reinforcement learning environment that enables the self-evolutionary training of clinically responsible and adaptive AI counselors.
@@ 1 Introduction
The intersection of Artificial Intelligence (AI) and mental healthcare has evolved from a theoretical curiosity into a burgeoning subfield of computational psychiatry [1, 2]. Historically, computer-mediated therapy was limited to rule-based systems like ELIZA, which relied on rigid pattern matching and lacked genuine semantic understanding. The advent of Large Language Models (LLMs) [3, 4, 5, 6] has fundamentally altered this landscape. Trained on internet-scale corpora and fine-tuned with reinforcement learning, modern LLMs demonstrate an emergent capability not only to parse complex syntax but to generate responses exhibiting empathy, reasoning, and context retention [7, 8]. This shift offers the potential to democratize psychological support through scalable, autonomous agents [9, 10].
@@
However, the transition from general-purpose chatbots to clinically reliable AI counselors introduces profound challenges. The therapeutic alliance—the collaborative bond between therapist and patient—relies on trust, confidentiality, and the ability to navigate ethical nuances. Generic LLMs, while fluent, are not inherently aligned with clinical protocols. They are prone to ”hallucinations” (fabricating medical facts), ”toxic positivity” (offering platitudes instead of validation), and safety failures in crisis scenarios. To mitigate these risks, recent research has introduced specialized models such as SoulChat [11], which optimizes for emotional resonance, and PsyLLM [10], which incorporates diagnostic reasoning via Chain-of-Thought (CoT) prompting. Furthermore, advanced frameworks like HealMe [12] and SMILE [13] have begun employing specialized architectural constraints and safety modules to ensure adherence to clinical guidelines.
@@
Despite these advancements, a critical gap persists between current research and real-world psychological counseling. First, real-world psychological counseling is not a static Question-Answering (QA) task, but a dynamic, longitudinal process. It requires a counselor to maintain memory continuity over multiple sessions, track evolving therapeutic goals, and navigate distinct stages—from initial case conceptualization to intervention and consolidation [14]. Second, effective counseling is rarely ”one-size-fits-all.” Human practitioners rely on diverse theoretical modalities (e.g., CBT, Psychoanalysis, Humanistic) and often adopt an integrative approach to suit the unique context of each client. Third, there is a stark deficiency in systematic evaluation methodologies. Traditional NLP metrics such as BLEU and ROUGE, while standard in text generation, are widely criticized in clinical contexts for their inability to capture therapeutic quality [15]. These metrics prioritize surface-level lexical overlap over semantic depth, failing to penalize ”over-validation” (indiscriminately agreeing with distorted thoughts) or measure the pragmatic appropriateness of a counselor’s intervention [16].
@@
In contrast, existing benchmarks for AI mental health primarily focus on single-turn dialogues, measuring generic empathy or safety in isolated contexts. Datasets such as PsyQA [17] and CPED provide valuable resources for advisory capabilities, yet they are predominantly composed of fragmented, single-session interactions or synthetic chat logs that lack the ”arc of therapy.” For example, most current datasets do not account for the temporal progression of a client’s symptoms or the gradual deepening of the therapeutic bond across multiple weeks. These static evaluations fail to capture the temporal progression of therapy (i.e., how a client evolves over time) and the theoretical versatility required to treat complex cases. An AI that performs well on a single-turn empathy benchmark may fail catastrophically when tasked with managing a long-term therapeutic alliance or adhering to a specific clinical framework.
@@
To bridge the gap between static benchmarking and dynamic clinical reality, we introduce PsychEval, a comprehensive, multi-session, and multi-therapy benchmark designed to rigorously assess AI counselors. Unlike previous datasets, PsychEval simulates the full trajectory of counseling. It incorporates a structured three-stage clinical flow—Case Conceptualization, Core Intervention, and Consolidation—spanning 6 to 10 sessions per client. To ensure theoretical depth, we construct a diverse dataset covering five major therapeutic modalities (Psychoanalytic, Behavioral, Cognitive-Behavioral, Humanistic-Existential, Postmodernist) and an Integrative approach, grounded in over 2,000 rigorous client profiles across six core psychological topics.
@@
  • RQ1: Can we train a highly realistic AI counselor? To ensure the model captures the longitudinal nature of therapy, we ground our training pipeline in high-fidelity, empirically sourced case reports rather than synthetic chat logs. Furthermore, by integrating a hierarchical skill taxonomy (mapping strategic meta-skills to executable atomic actions), we provide the necessary cognitive scaffolding that enables the model to sustain logical coherence and precise goal-tracking across complex, multi-session trajectories.
  • RQ2: How to train a multi-school AI counselor? To enable the mastery of distinct therapeutic orientations, we constructed a comprehensive dataset encompassing five specific therapies alongside an integrative category, covering a diverse array of counseling topics. Crucially, we employed a unified three-stage clinical framework (Case Conceptualization, Core Intervention, Consolidation) to model these varied data sources. This methodological design ensures that the agent can flexibly adapt to specific theoretical demands while adhering to a consistent, standardized structural progression across all counseling sessions.
  • RQ3: How to systematically evaluate an AI counselor? To move assessment beyond generic perplexity metrics, we establish a holistic evaluation framework operationalized across two key dimensions: Client-Level simulation fidelity and Counselor-Level clinical proficiency. Crucially, within these dimensions, we design a hybrid metric set comprising both therapy-specific indicators (to measure therapy distinctiveness) and shared indicators (to measure general clinical quality), ensuring a rigorous standard for both versatility and specialization.
Finally, we construct and open-source PsychEval. Comparative analysis with existing benchmarks demonstrates that PsychEval features the highest session depth and the closest alignment with real-world counseling dynamics to date. Furthermore, experimental evaluations confirm that the quality of our generated dataset significantly surpasses that of existing baselines. Crucially, PsychEval extends beyond a static evaluation testbed to serve as a high-fidelity reinforcement learning environment. By simulating diverse and evolving client states, it facilitates the self-evolutionary training of AI agents, providing a foundational ecosystem for developing clinically responsible and continuously improving AI counselors.

### v2

# PsychEval: A Multi-Session and Multi-Therapy Benchmark for High-Realism AI Psychological Counselor

@@ Abstract
To develop a reliable AI for psychological assessment, we introduce PsychEval, a multi-session, multi-therapy, and highly realistic benchmark designed to address three key challenges: 1) Can we train a highly realistic AI counselor? Realistic counseling is a longitudinal task requiring sustained memory and dynamic goal tracking. We propose a multi-session benchmark (spanning 6-10 sessions across three distinct stages) that demands critical capabilities such as memory continuity, adaptive reasoning, and longitudinal planning. The dataset is annotated with extensive professional skills, comprising over 677 meta-skills and 4577 atomic skills. 2) How to train a multi-therapy AI counselor? While existing models often focus on a single therapy, complex cases frequently require flexible strategies among various therapies. We construct a diverse dataset covering five therapeutic modalities alongside an integrative therapy with a unified three-stage clinical framework across six core psychological topics. 3) How to systematically evaluate an AI counselor? We establish a holistic evaluation framework with 18 therapy-specific and therapy-shared metrics across Client-Level and Counselor-Level dimensions. To We also construct over 2,000 diverse client profiles. Extensive experimental analysis fully validates the superior quality and clinical fidelity of our dataset. Crucially, PsychEval transcends static benchmarking to serve as a high-fidelity reinforcement learning environment that enables the self-evolutionary training of clinically responsible and adaptive AI counselors.
@@ 1 Introduction
The intersection of Artificial Intelligence and mental healthcare has evolved from a theoretical curiosity into a burgeoning subfield of computational psychiatry [1, 2]. Historically, computer-mediated therapy was confined to rule-based systems like ELIZA, which relied on rigid pattern matching and lacked genuine semantic understanding. The advent of Large Language Models (LLMs) [3, 4] has fundamentally transformed this landscape: trained on internet-scale corpora and fine-tuned via reinforcement learning, modern LLMs exhibit emergent capabilities not only to parse complex syntax but also to generate responses demonstrating empathy, reasoning, and context retention [5, 6]. This shift enables the potential to democratize psychological support through scalable, autonomous agents [7, 8].
@@
However, transitioning from general-purpose chatbots to clinically reliable AI counselors poses significant challenges. Therapeutic alliance, which depends on trust, confidentiality, and ethical navigation, requires more than fluency. Generic LLMs often lack clinical alignment, risking hallucinations, toxic positivity, and critical safety failures in crisis situations. To mitigate these risks, recent research has introduced specialized models such as SoulChat [9], which optimizes for emotional resonance, and PsyLLM [8], which incorporates diagnostic reasoning. Furthermore, advanced frameworks like HealMe [10] and SMILE [11] employ specialized architectural constraints and safety modules to ensure adherence to clinical guidelines.
@@
Despite these advances, a critical gap remains between current research and real-world psychological counseling. First, real counseling is not a static QA task but a dynamic, longitudinal process that requires memory continuity across sessions, tracking of evolving goals, and progression through distinct stages—from case conceptualization to intervention and consolidation [12]. Second, effective counseling is rarely one-size-fits-all; clinicians draw on diverse theoretical modalities (e.g., CBT, Psychoanalysis, Humanistic) and often integrate approaches tailored to each client. Third, systematic evaluation methods are lacking. Standard NLP metrics like BLEU and ROUGE are widely criticized in clinical contexts for capturing only surface lexical overlap, not therapeutic quality [13]. They fail to penalize “over-validation” (indiscriminate agreement with distorted thoughts) or assess the pragmatic appropriateness of interventions [14].
@@
In contrast, existing AI mental health benchmarks primarily assess single-turn dialogues, measuring generic empathy or safety in isolation. Datasets like PsyQA [15] offer valuable resources for advisory capabilities but consist largely of fragmented, single-session interactions or synthetic chat logs lacking the “arc of therapy.” For instance, they typically ignore the temporal progression of a client’s symptoms and the gradual deepening of the therapeutic bond over weeks. Such static evaluations fail to capture therapy’s temporal dynamics (i.e., how clients evolve) and the theoretical versatility needed for complex cases. Consequently, an AI excelling on single-turn empathy benchmarks may fail catastrophically when managing long-term therapeutic alliances or adhering to specific clinical frameworks.
@@
To bridge the gap between static benchmarking and dynamic clinical reality, we introduce PsychEval, a comprehensive, multi-session, and multi-therapy benchmark designed to rigorously assess AI counselors. Unlike previous datasets, PsychEval simulates the full trajectory of counseling. It incorporates a structured three-stage clinical flow: Case Conceptualization, Core Intervention, and Consolidation, spanning 6 to 10 sessions per client. To ensure theoretical depth, we construct a diverse dataset covering five major therapeutic modalities and an Integrative approach, grounded in over 2,000 rigorous client profiles across six core psychological topics.
@@
RQ1: Can we train a highly realistic AI counselor? To capture therapy’s longitudinal nature, we ground training in high-fidelity, empirically sourced case reports rather than synthetic chat logs and integrate a hierarchical skill taxonomy that maps strategic meta-skills to atomic actions, providing cognitive scaffolding for sustained logical coherence and precise goal-tracking across multi-session trajectories.
RQ2: How to train a multi-school AI counselor? To enable the mastery of distinct therapeutic orientations, we construct a comprehensive dataset encompassing five specific therapies alongside an integrative category, covering a diverse array of counseling topics. Crucially, we employ a unified three-stage clinical framework to model these varied data sources. This methodological design ensures that the agent can flexibly adapt to specific theoretical demands while adhering to a consistent structural progression across all sessions.
RQ3: How to systematically evaluate an AI counselor? To move assessment beyond generic perplexity metrics, we establish a holistic evaluation framework operationalized across two key dimensions: Client-Level simulation fidelity and Counselor-Level clinical proficiency. Crucially, within these dimensions, we design a hybrid metric set comprising both therapy-specific indicators and shared indicators, ensuring a rigorous standard for both versatility and specialization.
Finally, we construct and release PsychEval. Comparative analysis with existing benchmarks shows that PsychEval achieves the highest quality, session depth, and closest alignment with real-world counseling dynamics to date. Crucially, it extends beyond a static evaluation testbed to serve as a high-fidelity reinforcement learning environment, simulating diverse and evolving client states with rich reward functions to enable self-evolutionary training of AI agents and provide a foundational ecosystem for developing clinically responsible, continuously improving AI counselors.
@@ 3.1 Different Psychological Therapies
Over the past century, psychological counseling has evolved from singular theoretical origins into a multifaceted discipline, shaped by diverging and converging philosophical and scientific paradigms that have given rise to distinct therapeutic approaches. In constructing our benchmark, we primarily incorporate the following five major therapies: Psychodynamic [32], Behavioral [33, 34], Cognitive-Behavioral [35, 36], Humanistic-Existential [37], and Postmodernist [38], alongside an Integrative therapy. The descriptions of these therapies are provided in Section A in the Appendix.
3.2 Unified Counseling Workflow
Although different therapeutic approaches vary in their theories, techniques, and focal points, they exhibit common structural differentiations as the client’s problem emerges and evolves. Generally, the current counseling process can be divided into three distinct phases: evaluation, intervention, and consolidation [39, 12]. To ensure reproducibility and scientific rigor, we introduce a generalized framework that designates three phases: Case Conceptualization, Core Intervention, and Consolidation and Prevention (See Fig. 1).
Stage 1: Case Conceptualization (Assessment & Formulation)
This foundational phase centers on the client and typically spans 1–3 sessions. Its goals are to establish a secure therapeutic alliance and conduct a comprehensive assessment via intake interviews, gathering background, history, and symptom data. Guided by a theoretically informed conception rooted in clinical experience and research, the counselor synthesizes this information into an individualized case formulation and treatment plan, a personalized hypothesis about the etiology of the client’s difficulties that shapes the intervention trajectory.
Stage 2: Core Intervention (Action & Monitoring)
This central phase implements the course of therapy over 3–5 sessions, focusing on facilitating change through therapy-specific techniques (e.g., cognitive restructuring in CBT). Crucially, ongoing therapy monitoring creates a continuous feedback loop: if outcomes are unsatisfactory, the process loops back to reassess rapport, collect new data, or refine the case formulation. Thus, intervention is not rigid but a dynamic cycle of hypothesis testing and adaptation to the client’s real-time progress.
Stage 3: Consolidation and Prevention (Evaluation & Feedback)
The final 1–3 sessions focus on concluding evaluation to solidify gains and prepare for termination. Counselor and client collaboratively review the therapeutic journey to confirm intervention success. This evaluation both helps the client generalize skills to daily life (supporting relapse prevention) and provides the counselor with feedback to refine their internal guiding conceptions. The ultimate aim is to empower the client to sustain well-being independently.
Figure 2: The Flow of Structured Case Extraction.
4 Benchmark Construction
To bridge the gap between current AI capabilities and the complexity of real-world clinical practice, we construct a high-fidelity, multi-session benchmark. This benchmark integrates diverse therapeutic therapies to simulate the timely, accurate, and genuine adjustments required in psychological assessment. We employ advanced LLMs (e.g., GPT-5) to synthesize and reconstruct high-quality dialogues derived from empirically grounded clinical case reports. The construction pipeline proceeds through three primary phases: data collection, structured case extraction, and multi-stage dialogue construction.
4.1 Data Collection
To guarantee data authenticity and reliability, we curate a corpus of clinical case reports sourced from authoritative psychology journals, such as Theory and Practice of Psychological Counseling and Psychological Monthly. In contrast to unstructured transcripts from online consultation platforms (e.g., Yidianling), these peer-reviewed reports offer verified clinical frameworks, distinct theoretical orientations, and detailed longitudinal trajectories. Crucially, they originate from documented, real-world counseling sessions, thereby ensuring high ecological validity. Following rigorous cleaning and classification, the final dataset comprises 369 case reports. The distribution encompasses five major therapeutic modalities: Cognitive Behavioral Therapy accounts for 148 reports, while Psychoanalysis, Humanistic, and Postmodernist approaches each contribute 50. Additionally, the dataset includes 43 reports for Behaviorism and 28 for Integrative Therapy.
4.2 Structured Case Extraction
While LLMs possess strong document analysis capabilities, directly instructing them to reconstruct full counseling sessions from raw reports often yields suboptimal results, characterized by hallucinations, loss of clinical nuance, and flattened therapeutic arcs lacking cross-session continuity. To mitigate this, we implement an intermediate structured extraction phase. We convert raw text into a standardized schema consisting of four modules: Client Profile, Therapeutic Plan, Dialogue Features, and Therapeutic Skills (Fig. 2). This structured approach ensures the LLM captures the full granularity of the clinical context.
Client Profile
The Profile module constructs a structured persona integrating four dimensions: 1) Static Traits (Appendix Fig. 5), encompassing demographics, medical history, and linguistic features; 2) Clinical Presentation (Appendix Fig. 6), specifying the presenting problem, core demands, and counseling topics; 3) Growth Experience (Appendix Fig. 7), highlighting worldview-shaping life events; and 4) Therapy-Specific Parameters, extracting therapy-unique elements like core beliefs (CBT) or target behaviors (Behaviorism) to ensure professional distinctiveness.
Figure 3: The skill-informed generative pipeline for multi-stage dialogue construction.
Therapeutic Plan
As a strategic roadmap, this module generates a three-phase multi-session plan via the Profile and original report. While allowing flexible session counts, it enforces two constraints: the Timeline Constraint ensures chronological progression without foreseeing future events, while the Completeness Constraint permits logical inference of missing details. Each session plan is defined by five fields: Theme, Persona Links (activated profile data), Case Material (concrete quotes or tasks), Rationale (clinical reasoning), and Psychoeducation Skills (concepts for the client to learn).
Dialogue Features
To capture the micro-level dynamics of real counseling, this module extracts verbatim dialogue segments from the reports to serve as few-shot exemplars. This ensures the generated text moves beyond generic tones to reflect specific linguistic habits and interaction patterns. This module defines the Context (e.g., homework assignment), provides Verbatim Excerpts of 3–5 turn exchanges from the original report, and includes Stylistic Analysis instructions regarding sentence construction, lexical choice, and turn-taking habits derived from the real-world samples.
Therapeutic Skills
To bridge abstract theory with concrete action, we construct a hierarchical taxonomy enabling coarse-to-fine reasoning. Organized as a tree structure, Meta Skills (Strategic/Branch nodes) define therapeutic intent, while Atomic Skills (Tactical/Leaf nodes) define specific executable behaviors. 1) Meta Skills derived from authoritative textbooks (e.g., CBT: Basics and Beyond) represent high-level competencies. The extraction pipeline utilizes GPT-5 to parse digitized texts into a structured format, where each entry includes a Skill ID, Name, Description, and Parent IDs to maintain hierarchical lineage. 2) Atomic Skills represent granular verbal techniques extracted from practical clinical guides and transcripts. Unlike broad strategies, these are mapped directly to Meta Skill nodes to ensure theoretical alignment. To support dynamic decision-making, the schema is enriched with When to Use (contextual timing) and Trigger (specific client cues) alongside standard identifiers.
4.3 Multi-stage Dialogue Construction
Following case extraction, we develop a Skill-Informed Generative Pipeline (Fig. 3) to ensure realism and longitudinal continuity. The process comprises three phases: Pre-Session Preparation, In-Session Execution, and Post-Session Consolidation. This structure ensures interventions are grounded in assessment and maintains consistency across the counseling arc.
Phase 1: Pre-Session Preparation (Skill Utilization)
To define the intervention space, we propose a Coarse-to-Fine Skill Retrieval mechanism. First, Stage-Constraint Retrieval filters Meta Skills relevant to the current Therapeutic Plan phase. Then, functioning as a selector, GPT-5 identifies pertinent Atomic Skills based on session goals to construct a precise skill candidate set, ensuring the model possesses the necessary strategic toolkit without navigating the entire taxonomy.
Phase 2: In-Session Execution (Reasoning-Driven Generation)
Simulating human cognition, we employ Chain-of-Thought (CoT) [40] to generate an Internal Reasoning Trace prior to each utterance. This trace models four components: 1) Assessment (evaluating client state/subtext); 2) State Tracking (monitoring the alliance); 3) Skill Selection (choosing the optimal Atomic Skill); and 4) Strategy Formulation (determining tone/pacing). Decoupling reasoning from response ensures every interaction remains strategically grounded in clinical theory.
Phase 3: Post-Session Consolidation (Memory and Planning)
Ensuring cross-session consistency, this dual-mechanism module handles: 1) Information Synthesis, compressing the session into a Clinical Summary and updating the Client Profile with new disclosures; and 2) Trajectory Refinement, evaluating efficacy against the Therapeutic Plan to generate the Next Session Plan. Through Progress Assessment, the system determines whether to advance phases or schedule reinforcement, explicitly defining objectives and target skills for the subsequent encounter.
5 Holistic Evaluation Framework
To objectively quantify clinical and technical efficacy beyond traditional lexical metrics (e.g., BLEU), which fail to capture therapeutic intentionality, we establish a Holistic Evaluation Framework. Adopting an external supervisory paradigm, this methodology operationalizes validated psychometric instruments for automated LLM-based assessment. It integrates surrogate supervision to distinguish between universal and therapy-specific competencies, while systematically quantifying longitudinal therapeutic progress through multi-session trajectory analysis.
5.1 Counselor-Level Evaluation
The counselor-level evaluation assesses clinical proficiency, focusing on the delivery of theoretically grounded and ethically sound interventions. Adopting an external supervisory paradigm, we evaluate the agent’s longitudinal coherence and adherence to therapy-specific technical protocols.
Therapy-shared Metrics
These metrics evaluate universal competencies using established instruments and specialized AI assessments. We utilize the Working Alliance Inventory (WAI; refer to Appendix Fig. 36 for the detailed prompt design), Helping Transaction Audit Inventory Scale (HTAIS) [41], and Real Relationship-Observer (RRO) [42] to measure core relational factors (alliance, integrity, genuineness). To facilitate a more holistic assessment of dataset quality, we propose a customized evaluation framework that rigorously examines four key domains: clinical perception, longitudinal strategy, therapeutic depth, and ethical safety adherence. The specific prompt designs for these dimensions are detailed in Fig. 39,41,40 et al.
Cognitive-Behavioral 148 7.5 22.0 119.9 371.7 10 103 1368
Psychodynamic 50 8.0 22.0 86.4 334.7 10 210 1144
Behavioral 43 7.2 28.0 70.3 306.0 9 156 854
Humanistic-Existential 50 7.7 25.1 81.6 326.7 10 136 752
Postmodernist 50 7.5 26.7 73.0 302.1 10 72 459
Integrative 28 7.7 25.9 75.4 308.3 10 - -
Total 369 7.6 24.1 92.9 336.2 10 677 4577
[…]
  Dataset Reasoning Evaluation MultiSess MultiTherapy Integrative MultiStage Skill OpenSource
CounselingBench [7]
PsyDTCorpus [43]
CACTUS [44]
PsyDial [45]
SimPsyDial [46]
Cpsycoun [47]
MindChat [48]
SMILE [11]
MusPsy [49]
Psy-Insight [50]
PsychEval (Our)
Table 2: Comparison with existing benchmarks on key characteristics. Reasoning denotes turn-level reasoning; Evaluation refers to a comprehensive assessment framework; Integrative indicates support for integrative therapy.
(per Case) (per Session) (per Session) (per Session)
Overall (Total) 7.58 4.65 5.73 7.58
Stage 1: Intake & Conceptualization 1.84 7.47 5.91 8.56
Stage 2: Core Intervention 4.13 3.98 5.92 7.51
Stage 3: Consolidation & Termination 1.62 3.14 5.03 6.64
Figure 4: Statistical information of skills and topics.
Therapy-specific Metrics
These metrics assess theoretical fidelity across orientations using specialized scales. We employ the Cognitive Therapy Rating Scale (CTRS) for CBT competencies (e.g., Socratic questioning) and the Psychodynamic Supervision Checklist (PSC) for insight into unconscious conflicts. Furthermore, the Therapist Empathy Scale (TES) and Therapeutic Focus Scale (EFT-TFS) measure empathic resonance and affective processing in Humanistic/Behavioral approaches, while the Motivational Interviewing Treatment Integrity (MITI) scale gauges change discourse evocation in Postmodern therapies.
5.2 Client-Level Evaluation
The Client-Level evaluation focuses on simulation fidelity, assessing whether simulated clients maintain clinical consistency and authentic psychological shifts. We employ an LLM-based supervisor to convert qualitative dialogue into quantifiable clinical data via standardized scoring.
Therapy-shared Metrics
These metrics assess internal states and dyadic qualities using four instruments. We utilize the Symptom Checklist-90 (SCL-90) to track phenotype consistency and symptom reduction, and the Positive and Negative Affect Schedule (PANAS) for acute affective transitions. Furthermore, the Real Relationship Inventory (RRO) evaluates counselor authenticity, while the Session Rating Scale (SRS) monitors the working alliance and potential relationship ruptures.
Therapy-specific Metrics
These metrics validate theoretical alignment using five specialized instruments. We employ the Beck Depression Inventory-II (BDI-II) to verify cognitive symptom reduction in CBT and the Inventory of Personality Organization (IPO) to track personality integration in Psychodynamic therapy. The State-Trait Anxiety Inventory (STAI) and Client-Centered Therapy Criteria (CCT) measure anxiety and self-acceptance in Behavioral and Humanistic approaches, respectively, while SFBT Indicators monitor solution-oriented discourse in Postmodern interventions.
6 Analysis of Benchmark
6.1 Statistic Information
Table 1 presents the statistical profile of PsychEval, a benchmark distinguished by unprecedented longitudinal depth and high-fidelity information density. It comprises 369 cases across five therapeutic genres, breaking the ”single-session” barrier of prior works with an average of 7.6 sessions per case (up to 10). This multi-session structure provides the temporal span essential for evaluating memory continuity and long-term planning. Interactions exhibit professional-grade density, averaging 24.1 turns per session and 94.7 words per turn, with realistic variation across genres: Cognitive-Behavioral Therapy has the highest volume (119.9 words/turn) due to its psychoeducational nature, while Behavioral approaches remain concise ($\sim$70 words/turn), reflecting authentic clinical practices.
Moreover, PsychEval sets a benchmark for interpretable counseling through a hierarchical taxonomy of 677 Meta and 4,577 Atomic skills (Fig. 4), enabling navigation from strategic planning to tactical intervention. As shown in Table 3, the framework’s structural integrity is validated by its alignment with a professional three-stage clinical model: metric distributions mirror authentic therapeutic dynamics, from high personalization in intake to sustained duration and material richness during core intervention.
PsyDTCorpu[43] 4760 12 1 1 44.8 18.1 4760 86054
CACTUS [44] 31577 9 1 1 26.4 15.6 31577 491316
PsyDial (D4) [45] 2382 60 1 3 32.4 37.8 2382 90033
SimPsyDial [46] 1000 60 1 3 62.5 12.9 1000 12948
Cpsycoun [47] 3084 9 1 7 39.0 8.0 3084 24808
SMILE [11] 55165 60 1 - 81.3 5.8 55165 318395
ESConv [22] 1300 12 1 - 16.3 14.5 1300 18864
MusPsy [49] 1400 - 6.2 1 27.0 28.6 - -
[…]
PsychEval (Our) 369 6 7.6 5 92.9 24.1 2798 67314
  Therapy-Shared Therapy-Specific
HTAIS RRO WAI Custom CTRS PSC TES EFT-TFS MITI
CACTUS [44] 4.92 7.12 6.08 3.64 1.41 - - - -
Cpsycoun [47] 4.74 6.70 5.43 3.94 1.30 2.87 2.87 1.04 2.94
PsyDial [45] 4.25 6.39 4.15 4.48 4.47 - 5.71 - -
PsyDTCorpus [43] 5.50 7.44 6.16 4.40 1.35 - - - -
SimPsyDial [46] 4.55 6.02 4.87 4.10 1.96 3.10 3.52 - -
SMILE [11] 5.32 7.07 5.37 4.13 - - - - -
[…]
ESConv [22] 4.60 6.79 4.66 4.19 - - 3.26 - -
PsychEval (Our) 6.40 7.64 7.26 7.36 9.19 7.27 7.36 3.14 5.88
  Therapy-Shared Therapy-Specific
[…]
CACTUS [44] 6.37 4.92 3.09 7.47 1.84 - - - -
Cpsycoun [47] 6.15 4.69 2.94 6.83 1.20 0.64 4.04 5.24 -
PsyDial [45] 6.36 4.12 3.23 6.99 2.33 - 6.36 - -
PsyDTCorpus [43] 6.94 4.83 3.06 8.01 1.79 - - - -
SimPsyDial [46] 6.11 4.72 2.97 7.63 1.63 1.27 4.85 - -
SMILE [11] 6.52 4.83 3.04 6.80 - - - - -
[…]
ESConv [22] 6.49 4.68 3.07 6.36 - - 4.66 - -
PsychEval (Our) 6.70 5.27 1.48 8.29 1.87 2.17 8.41 7.84 5.57
  Therapy-Shared Therapy-Specific
[…]
PsychEval (Our) 0.07 0.17 -0.06 0.08 -0.45 -0.09 0.14 0.16 0.28
6.2 Comparison with Existing Benchmark
To contextualize PsychEval’s contribution, we compare it multidimensionally against representative datasets. As shown in Tables 4 and 2, PsychEval shifts focus from large-scale but shallow interactions to high-fidelity, longitudinal, and skill-aware assessment.
First, it sets a new standard in longitudinal depth and information density. While most existing datasets are limited to single-session interactions and low verbal complexity (e.g., ESConv: 16.3 words/turn), PsychEval provides an unprecedented average of 7.6 sessions per case and 92.9 words per turn. This extended temporal span, coupled with rigorous “chit-chat” filtering, captures the evolving dynamics of real-world therapy rather than static snapshots. Second, PsychEval is the most functionally complete benchmark to date. As Table 2 shows, it is the only open-source dataset that simultaneously supports turn-level reasoning, multi-stage clinical workflows, and hierarchical skill annotations. Unlike single-modality or supervision-lacking datasets, it integrates five distinct therapeutic genres plus an integrative approach. This combination of longitudinal fidelity and structured expert annotations advances the field beyond basic dialogue modeling toward skill-aware, planning-capable AI counselors.
6.3 Data Quality Assessment
To rigorously validate the clinical fidelity of PsychEval, we conducted a comparative quality assessment against eight established benchmarks over psychological metrics. As shown in Tables 5, 6, and 7, PsychEval demonstrates superior performance across both counselor and client dimensions.
In Table 5, PsychEval excels in both Therapy-Shared and Therapy-Specific dimensions, outperforming datasets such as PsyDTCorpus and Cpsycoun. It achieves SOTA scores in Working Alliance (WAI: 7.26) and Helper Skills (HTAIS: 6.40), and shows especially strong therapy-specific adherence, with CTRS at 9.19 and PSC at 7.27, nearly double prior models like PsyDial and SimPsyDial. This confirms that PsychEval captures both empathetic tone and precise clinical interventions.
For client realism and therapeutic effectiveness, we evaluate post-session psychological states and longitudinal changes. Table 6 shows PsychEval clients report the highest satisfaction and alliance (SRS: 8.29) and the largest reduction in symptom distress (SCL-90: 1.48). Crucially, authentic therapy is defined by change over time. Table 7 presents adjacent differences (session $t+1$ minus session $t$) as proxies for progress: PsychEval exhibits a coherent “healing” trajectory—symptom measures (BDI-II, IPO) decrease significantly (-0.45, -0.09), while positive capacities (CCT, SFBT) increase (+0.14, +0.16). In contrast, Psy-Insight shows counter-therapeutic trends: IPO rises (+0.04), and CCT/SFBT decline (-0.19, -0.13). Thus, PsychEval is the only benchmark that authentically models real-world psychological evolution and positive transformation.
7 Conclusions and Future Work
We introduce PsychEval, a benchmark and training ecosystem that aligns AI capabilities with professional psychological assessment demands. Built from 369 authentic clinical case reports, it captures therapy’s longitudinal complexity with an average of 7.6 sessions per case. PsychEval combines a three-stage clinical framework with a hierarchical taxonomy of over 4,500 atomic skills, enabling AI agents to perform coarse-to-fine clinical reasoning. It significantly outperforms existing benchmarks, achieving a CTRS score of 9.19 and a WAI score of 7.26. Beyond evaluation, PsychEval serves as a high-fidelity reinforcement learning environment for training clinically responsible, logically coherent, and ethically aligned AI counselors. Future work will use PsychEval to develop a self-evolving AI counselor for autonomous, adaptive mental health care.
Appendix A Different Psychological Therapies